Suppose X1,X2,...,Xn is a RS from Bernoulli distribution. Define the estimator theta hat=1/n+√n( summation Xi+√n/n) Show that theta hat is a consistent estimator of theta



1
Expert's answer
2022-01-10T16:23:28-0500

estimator of Bernoulli distribution:

"\theta=p=\frac{\sum x_i}{n}"


"\hat{\theta }" is a consistent estimator if

"\displaystyle \lim_{n\to \infin} Pr(|\hat{\theta}-\theta|> \varepsilon)=0"

for all "\varepsilon >0"


then:

"\hat{\theta}-\theta=\frac{1}{n+\sqrt n}( \sum x_i+\sqrt n)/n-\frac{\sum x_i}{n}"


"\displaystyle \lim_{n\to \infin}|\hat{\theta}-\theta|=0"


"\displaystyle \lim_{n\to \infin} Pr(0> \varepsilon)=0"


so, "\hat{\theta }" is a consistent estimator of "\theta"